4. v defined by vn = n! + 2, n ≥ 1.

(a) Find v3

(b) Find  Σ 4 on top , i=1 at the bottom and Vi on the right hand side of Sigma

(c) Is v increasing, decreasing, non-increasing or non-decreasing?


1
Expert's answer
2021-11-01T19:54:58-0400

(a)


"v_3=3!+2=8"

(b)


"\displaystyle\sum_{i=1}^{4}v_i=v_1+v_2+v_3+v_4"

"=1!+2+2!+2+3!+2+4!+2=41"

(c)


"v_{n+1}=(n+1)!+2>n!+2=v_n, n\geq 1"

The sequence "\{v_n\}" is (strictly) increasing.